# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

### v2

@@ III X17 couplings to charm and strange quarks
The VMD assumption has been phenomenologically successful in explaining TFF behaviors in similar decays. The effective pole mass refers to the mass of the vector meson resonance that is close to the energy scale of the decaying particle. In the analysis of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\Lambda_{\psi(2S)\eta_{c}}$ corresponds to the mass of a higher excited state, $m_{\psi(3770)}=(3773.7\pm 0.4)\,\mathrm{MeV/c^{2}}$ Gu:2019qwo . This framework has provided a coherent explanation for the experimental observations in the Charmonium decays. From Eq. (7), the ratio $R_{\psi(2S)}^{X}$ between the decay rate of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$ and that of $\psi(2S)\rightarrow\eta_{c}\gamma$ is given by
$R_{\psi(2S)}^{X}=9\varepsilon_{c}^{2}/4.$ (18)
This is an approximated result in the limit $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\psi(2S)\eta_{c}\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. The measurement by BESIII gives ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}\right)=$ $\left(3.77\pm 0.40_{\text{stat }}\pm 0.18_{\text{syst }}\right)\times 10^{-5}$ BESIII:2022jde . Utilizing the branching ratio ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}\gamma\right)=(3.6\pm 0.5)\times 10^{-3}$ ParticleDataGroup:2024cfk , one obtains the ratio $R^{\rm exp}_{\psi(2S)}=(1.1\pm 0.3)\times 10^{-2}$ as reported in BESIII:2022jde . The theoretical prediction for the photon-mediated decay rate given by the VMD model is $R_{\psi(2S)}^{\gamma}=8.9\times 10^{-3}$. The tiny difference between $R_{\psi(2S)}^{\gamma}$ and $R^{\rm exp}_{\psi(2S)}$ strongly constrain $R_{\psi(2S)}^{X}$, i.e., the coupling parameter $\varepsilon_{c}$.
The decay $\phi\rightarrow\eta e^{+}e^{-}$is a transition from a pure $\bar{s}s$ state to a mixed state involving up ($\bar{u}u$), down ($\bar{d}d$), and strange ($\bar{s}s$) quark pairs, emitting an electron-positron pair in the process. The measurements by SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk yield the branching ratio for $\phi\rightarrow\eta e^{+}e^{-}$ as $\left(1.19\pm 0.19_{\text{stat. }}\pm\right.$ $\left.0.07_{\text{syst. }}\right)\times 10^{-4},\left(1.14\pm 0.10_{\text{stat. }}\pm 0.06_{\text{syst. }}\right)\times 10^{-4}$ and $\left(1.075\pm 0.007_{\text{stat. }}\pm 0.038_{\text{syst. }}\right)\times 10^{-4}$, respectively. The VMD approach is utilized to estimate the lepton mass spectrum for the electromagnetic (EM) Dalitz decay $\phi\rightarrow\eta e^{+}e^{-}$. Following the similar approach as that of Eq. (15), we obtain the relation
@@
[…]
$R_{\phi}^{X}=9\varepsilon_{s}^{2}.$ (21)
Similar to the case of $\psi(2S)$ decay, we obtain the above relation with the approximations $m_{e}^{2}/q^{2}\equiv m_{e}^{2}/m_{X}^{2}\to 0$, and $\mathcal{F}_{\phi\eta\gamma}\left(q^{2}=m_{X}^{2}\right)\to 1$. Given the measurements on $R_{\phi}$ from three separate experiments and the theoretical prediction, $R_{\phi}^{\gamma}=8.30\times 10^{-3}$, we are able to determine the allowed range for $R_{\phi}^{X}$.
@@
Figure 3: The allowed ranges for (a) $\left(\varepsilon_{c},\varepsilon_{s}\right)$ and (b) $\left(\varepsilon_{c},\varepsilon_{u}\right)$ extracted by the data of $D^{*+}_{s}$, $D^{*0}$, $\psi(2S)$ and $\phi$ decays. The values of $\varepsilon_{u}$ extracted by Refs. Feng:2016jff ; Fornal:2017msy and Ref. Denton:2023gat are also shown.
The $1\sigma$ parameter range allowed by the ratio between $\Gamma(\psi(2S)\to\eta_{c}e^{+}e^{-})$ BESIII:2022jde and $\Gamma(\psi(2S)\to\eta_{c}\gamma)$ is denoted by the pink area in Fig. 3. The left panel depicts the allowed range for the $\left(\varepsilon_{c},\varepsilon_{s}\right)$ pair while the right panel depicts that for $\left(\varepsilon_{c},\varepsilon_{u}\right)$. Furthermore, the left panel of Fig. 3 also shows $1\sigma$ ranges for $\varepsilon_{s}$ from the data of SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk on $\phi\to\eta e^{+}e^{-}$, denoted by yellow, gray, and purple areas, respectively. The blue band is the allowed range for $\left(\varepsilon_{c},\varepsilon_{s}\right)$ by the data of $D^{*+}_{s}\to D^{+}_{s}e^{+}e^{-}$ CLEO:2011mla . On the right panel of Fig. 3, the $1\sigma$ allowed range for $\left(\varepsilon_{c},\varepsilon_{u}\right)$ by the data of $D^{*0}\to D^{0}e^{+}e^{-}$ is denoted by the green areas. The parameter ranges for various $D^{*}$ meson decays are determined with $\lambda=-0.289\ {\rm GeV}^{-1}$. Finally the extracted values for $\varepsilon_{u}$ from Refs. Feng:2016jff ; Fornal:2017msy and Ref. Denton:2023gat are denoted by black dashed curves and red band, respectively.
Table 2: The measured values and the corresponding Standard Model (SM) predictions of four decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons, respectively.
@@
$\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ $(11.08\pm 0.9)\times 10^{-3}$ $6.7\times 10^{-3}$
$\Gamma(D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-})/\Gamma(D_{s}^{*+}\rightarrow D_{s}^{+}\gamma)$ $\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ $6.7\times 10^{-3}$
$\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ $(11\pm 3)\times 10^{-3}$ $8.9\times 10^{-3}$
$\Gamma(\phi\rightarrow\eta e^{+}e^{-})/\Gamma(\phi\rightarrow\eta\gamma)$ $(8.31\pm 0.33)\times 10^{-3}$ $8.30\times 10^{-3}$
We employ the $\chi^{2}$ statistical test for fitting the coupling parameters, $\varepsilon_{c},\varepsilon_{s}$, and $\varepsilon_{u}$ with the four measurements given in Table 2. For $\phi\to\eta e^{+}e^{-}$ decay, we only employ the data by KLOE-2 KLOE-2:2014hfk since its uncertainty is much smaller than those of earlier measurements. We minimize the $\chi^{2}$ defined as
$\chi^{2}=\sum_{i=1}^{4}\frac{\left(R_{i}^{\rm th}\left(\varepsilon_{c},\varepsilon_{s},\varepsilon_{u},\lambda\right)-R_{i}^{\rm ob}\right)^{2}}{\sigma_{i}^{2}}+\frac{\left(\lambda-\lambda_{\rm best\,fit}\right)^{2}}{\sigma_{\lambda}^{2}},$ (22)
where $R_{i}^{\rm th}$ denotes the theoretically predicted value for the $i$-th decay ratio, which depends on the parameters $\varepsilon_{c}$, $\varepsilon_{s}$, $\varepsilon_{u}$, and $\lambda$ (for $i=D^{*0},\ D^{*+}_{s}$). $R_{i}^{\rm ob}$ refers to the observed value for the $i$-th decay ratio with $\sigma_{i}$ the corresponding uncertainty for the measurement. The parameter $\lambda$ is defined in Eq. (11) with its range given by Eq. (13). We hence take $\lambda_{\rm best\,fit}=-0.289$ GeV-1 and $\sigma_{\lambda}=0.016$ GeV-1. Our fitting uses the measured decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons as outlined in Table 2.

### v3

@@ III X17 couplings to charm and strange quarks
The VMD assumption has been phenomenologically successful in explaining TFF behaviors in similar decays. The effective pole mass refers to the mass of the vector meson resonance that is close to the energy scale of the decaying particle. In the analysis of $\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}$, $\Lambda_{\psi(2S)\eta_{c}}$ corresponds to the mass of a higher excited state, $m_{\psi(3770)}=(3773.7\pm 0.4)\,\mathrm{MeV/c^{2}}$ Gu:2019qwo . This framework has provided a coherent explanation for the experimental observations in the Charmonium decays. The ratio $R_{\psi(2S)}^{X}$ between the decay rate of $\psi(2S)\rightarrow\eta_{c}X\rightarrow\eta_{c}e^{+}e^{-}$ and that of $\psi(2S)\rightarrow\eta_{c}\gamma$ is given by
$R_{\psi(2S)}^{X}\equiv\frac{\Gamma^{X}(\psi(2S)\to\eta_{c}e^{+}e^{-})}{\Gamma(\psi(2S)\to\eta_{c}\gamma)}\approx\frac{\Gamma(\psi(2S)\to\eta_{c}X)\times{\rm BR}(X\to e^{+}e^{-})}{\Gamma(\psi(2S)\to\eta_{c}\gamma)},$ (19)
where the narrow width approximation for the X17 boson propagator has been used. With ${\rm BR}(X\to e^{+}e^{-})=1$, we have $R^{X}_{\psi(2S)}=\Gamma(\psi(2S)\to\eta_{c}X)/\Gamma(\psi(2S)\to\eta_{c}\gamma)$. Applying Eq. (3), we arrive at
$R_{\psi(2S)}^{X}=\frac{F^{2}_{\psi(2S)\eta_{c}X}(m_{X}^{2})}{F^{2}_{\psi(2S)\eta_{c}\gamma}(0)}\times\frac{\lambda^{1/2}(m_{\psi(2S)}^{2},m_{\eta_{c}}^{2},m_{X}^{2})}{\lambda^{1/2}(m_{\psi(2S)}^{2},m_{\eta_{c}}^{2},0)}.$ (20)
From Eq. (15), we have $F^{2}_{\psi(2S)\eta_{c}X}(m_{X}^{2})/F^{2}_{\psi(2S)\eta_{c}\gamma}(0)=\mathcal{F}^{2}_{\psi(2S)\eta_{c}X}(m_{X}^{2})$. Applying Eqs. (17) and (18), we obtain $\mathcal{F}^{2}_{\psi(2S)\eta_{c}X}(m_{X}^{2})=(9\varepsilon_{c}^{2}/4)\times(1-m_{X}^{2}/\Lambda_{\psi(2S)\eta_{c}}^{2})^{-2}$, which approaches to $9\varepsilon_{c}^{2}/4$ as $m_{X}^{2}/\Lambda_{\psi(2S)\eta_{c}}^{2}\to 0$. Furthermore, $\lambda^{1/2}(m_{\psi(2S)}^{2},m_{\eta_{c}}^{2},m_{X}^{2})/\lambda^{1/2}(m_{\psi(2S)}^{2},m_{\eta_{c}}^{2},0)\to 1$ due to $m_{X}^{2}\ll m_{\psi(2S)}^{2}\ (m_{\eta_{c}}^{2})$. We conclude that $R_{\psi(2S)}^{X}\to 9\varepsilon_{c}^{2}/4$.
The measurement by BESIII gives ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-}\right)=$ $\left(3.77\pm 0.40_{\text{stat }}\pm 0.18_{\text{syst }}\right)\times 10^{-5}$ BESIII:2022jde . Utilizing the branching ratio ${\rm BR}\left(\psi(2S)\rightarrow\eta_{c}\gamma\right)=(3.6\pm 0.5)\times 10^{-3}$ ParticleDataGroup:2024cfk , one obtains the ratio $R^{\rm exp}_{\psi(2S)}=(1.1\pm 0.3)\times 10^{-2}$ as reported in BESIII:2022jde . The theoretical prediction for the photon-mediated decay rate given by the VMD model is $R_{\psi(2S)}^{\gamma}=8.9\times 10^{-3}$. The tiny difference between $R_{\psi(2S)}^{\gamma}$ and $R^{\rm exp}_{\psi(2S)}$ strongly constrain $R_{\psi(2S)}^{X}$, i.e., the coupling parameter $\varepsilon_{c}$.
[…]
@@
For the electromagnetic decay process $\phi\rightarrow\eta e^{+}e^{-}$, the pole mass in the VMD would naturally be a $s\bar{s}$ resonance that is close to the decaying $\phi$ meson in mass. Since $\phi(1680)$ is a higher-mass excited state of the $\phi$ meson, one thus takes $\Lambda_{\phi\eta}=m_{\phi(1680)}=(1680\pm 20)\,\mathrm{MeV}/\mathrm{c}^{2}$. In the scenario of $\mathrm{X}17$ boson mediated $\phi\rightarrow\eta e^{+}e^{-}$ decay, i.e., $\phi\to\eta X$ with $X\to e^{+}e^{-}$, we obtain
$R_{\phi}^{X}=\frac{F^{2}_{\phi\eta X}(m_{X}^{2})}{F^{2}_{\phi\eta\gamma}(0)}\times\frac{\lambda^{1/2}(m_{\phi}^{2},m_{\eta}^{2},m_{X}^{2})}{\lambda^{1/2}(m_{\phi}^{2},m_{\eta}^{2},0)}\to 9\varepsilon_{s}^{2}.$ (23)
The above limit arises from $F^{2}_{\phi\eta X}(m_{X}^{2})/F^{2}_{\phi\eta\gamma}(0)\to 9\varepsilon_{s}^{2}$ due to Eqs. (15), (21) and (22), and $\lambda^{1/2}(m_{\phi}^{2},m_{\eta}^{2},m_{X}^{2})/\lambda^{1/2}(m_{\phi}^{2},m_{\eta}^{2},0)\to 1$ because of $m_{X}^{2}\ll m_{\phi}^{2}\ (m_{\eta}^{2})$. Given the measurements on $R_{\phi}$ from three separate experiments and the theoretical prediction, $R_{\phi}^{\gamma}=8.30\times 10^{-3}$, we are able to determine the allowed range for $R_{\phi}^{X}$.
@@
Figure 3: The allowed ranges for (a) $\left(\varepsilon_{c},\varepsilon_{s}\right)$ and (b) $\left(\varepsilon_{c},\varepsilon_{u}\right)$ extracted by the data of $D^{*+}_{s}$, $D^{*0}$, $\psi(2S)$ and $\phi$ decays. The values of $\varepsilon_{u}$ extracted by Refs. Feng:2016jff ; Fornal:2017msy (black dashed lines) and Ref. Denton:2023gat (red band) are also shown.
The $1\sigma$ parameter range allowed by the ratio between $\Gamma(\psi(2S)\to\eta_{c}e^{+}e^{-})$ BESIII:2022jde and $\Gamma(\psi(2S)\to\eta_{c}\gamma)$ is denoted by the pink area in Fig. 3. The left panel depicts the allowed range for the $\left(\varepsilon_{c},\varepsilon_{s}\right)$ pair while the right panel depicts that for $\left(\varepsilon_{c},\varepsilon_{u}\right)$. Furthermore, the left panel of Fig. 3 also shows $1\sigma$ ranges for $\varepsilon_{s}$ from the data of SND Achasov:2000ne , CMD-2 CMD-2:2000qla and KLOE-2 KLOE-2:2014hfk on $\phi\to\eta e^{+}e^{-}$, denoted by yellow, gray, and purple areas, respectively. The blue band is the allowed range for $\left(\varepsilon_{c},\varepsilon_{s}\right)$ by the data of $D^{*+}_{s}\to D^{+}_{s}e^{+}e^{-}$ CLEO:2011mla . On the right panel of Fig. 3, the $1\sigma$ allowed range for $\left(\varepsilon_{c},\varepsilon_{u}\right)$ by the data of $D^{*0}\to D^{0}e^{+}e^{-}$ is denoted by the green areas. The parameter ranges for various $D^{*}$ meson decays are determined with $\lambda=-0.289\ {\rm GeV}^{-1}$. Finally, the extracted values for $\varepsilon_{u}$ from Refs. Feng:2016jff ; Fornal:2017msy and Ref. Denton:2023gat are denoted by black dashed curves and red band, respectively.
Table 2: The measured values and the corresponding SM predictions of four decay ratios of $D^{*0}$, $D_{s}^{*+}$, $\psi(2S)$, and $\phi$ mesons, respectively.
@@
$\Gamma(D^{*0}\rightarrow D^{0}e^{+}e^{-})/\Gamma(D^{*0}\rightarrow D^{0}\gamma)$ $(11.08\pm 0.90)\times 10^{-3}$ BESIII:2021vyq $6.4\times 10^{-3}$
$\Gamma(D_{s}^{*+}\rightarrow D_{s}^{+}e^{+}e^{-})/\Gamma(D_{s}^{*+}\rightarrow D_{s}^{+}\gamma)$ $\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla $6.5\times 10^{-3}$
$\Gamma(\psi(2S)\rightarrow\eta_{c}e^{+}e^{-})/\Gamma(\psi(2S)\rightarrow\eta_{c}\gamma)$ $(11\pm 3)\times 10^{-3}$ BESIII:2022jde $8.9\times 10^{-3}$
$\Gamma(\phi\rightarrow\eta e^{+}e^{-})/\Gamma(\phi\rightarrow\eta\gamma)$ $(8.31\pm 0.33)\times 10^{-3}$ KLOE-2:2014hfk $8.30\times 10^{-3}$
Figure 4: The distribution ${\rm d}R_{D^{*0}}/{\rm d}q^{2}$ as a function of $e^{+}e^{-}$ invariant mass $q^{2}$ with the bin width $\Delta q^{2}=200\ {\rm MeV}^{2}$. The SM contribution, ${\rm d}R^{\gamma}_{D^{*0}}/{\rm d}q^{2}$, and X17 contribution, ${\rm d}R^{X}_{D^{*0}}/{\rm d}q^{2}$, with $200\ {\rm MeV}^{2}$ invariant-mass binning are denoted in pink and blue, respectively. For comparisons, the analytic result for SM contribution ${\rm d}R^{\gamma}_{D^{*0}}/{\rm d}q^{2}$ is represented by the red curve.
As summarized in Table 2, the measured value of $R_{D^{*0}}$ is much greater than the SM predicted value. The difference between the two is $\delta R_{D^{*0}}=(4.68\pm 0.90)\times 10^{-3}$, which is roughly $3/4$ of $R^{\gamma}_{D^{*0}}$. If such a difference is due to X17 contribution, there should exist a peak in the distribution ${\rm d}R_{D^{*0}}/{\rm d}q^{2}$ at $q^{2}=m_{X}^{2}$. In Fig. 4 we plot ${\rm d}R_{D^{*0}}/{\rm d}q^{2}$ with the energy binning adopted by Ref. BESIII:2021vyq , i.e., $\Delta q^{2}=200\ {\rm MeV}^{2}$. The SM contribution, ${\rm d}R^{\gamma}_{D^{*0}}/{\rm d}q^{2}$, and X17 contribution, ${\rm d}R^{X}_{D^{*0}}/{\rm d}q^{2}$, with the $200\ {\rm MeV}^{2}$ invariant-mass binning are denoted in pink and blue, respectively. For comparisons, the analytic result for SM contribution ${\rm d}R^{\gamma}_{D^{*0}}/{\rm d}q^{2}$ is represented by the red curve. Since $m_{X}^{2}\approx 280\ {\rm MeV}^{2}$, X17 contribution to ${\rm d}R_{D^{*0}}/{\rm d}q^{2}$ is at the second bin. As a result, the second bin would surpass the first bin in events. This is in contrary to the case with ${\rm d}R^{\gamma}_{D^{*0}}/{\rm d}q^{2}$ where event number in the first bin dominates that of any other bin. Although the spectral plot in BESIII:2021vyq does not show this feature, it remains inconclusive since it is also stated in that paper that the $q^{2}$ distribution measurement is not yet meaningful due to the limited statistics.